A particle of mass 'm' moves along a circle of radius R with a normal acceleration varying with time as a=b*t^2 where b is a constant.The time dependence of the power developed by all the forces acting on the particle is:a) 2mRbt b)mRbt c)4mRbt d)3mRat

Grade:12

3 Answers

Ramesh V
14 years ago

normal accln: an = v2/R = bt2

so v2 = bt2R

Kinetic Energy(E) = 1/2 mv2 = 1/2*mbt2R

power P =dE/dt = d(1/2*mbt2R) / dt

                            P =mbt2R

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and
we will get you the answer and detailed solution very quickly.We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch
 

JAGABANDHU PANDA
14 years ago

Hi

I would like to make some corrections to the reply posted by Mr Ramesh. I think after differentiation of "E" we would be getting P = mRbt which is the answer i. e. the answer is the option (b).

JAGABANDHU PANDA
14 years ago

Hi

The differentiation needs to be corrected and accordingly the answer is (b)

JAGABANDHU PANDA

IIT Delhi